OpenStudy (anonymous):

Let's say that the variable 'x' is definitely some negative number. So if I wanted to solve: \[x^2=4\] I get: \[\pm \sqrt{x^2} = \pm \sqrt{4} \\ \pm x = \pm 2\] I would have to take the positive value of 'x' and the negative value of '2' to make this true...is it okay to only take a positive square root of one side of the equation and the negative square root of the other?

OpenStudy (anonymous):

\[\sqrt{x ^{2}}=\left| x \right|\]

OpenStudy (anonymous):

it would be \[x=\pm2\] as per me

OpenStudy (anonymous):

So every time I write it out, I have to put \[\pm \sqrt{x^2} = \pm \sqrt{4} \\ \pm |x| = \pm2 \\ -|x| = -2\]? And I have to solve the absolute value equation every time?

OpenStudy (anonymous):

why \[\pm \sqrt{x ^{2}}???\]

OpenStudy (anonymous):

Because when you take the square root of a number, you don't know if 'x' is positive or negative

OpenStudy (anonymous):

so what......that doesnt mean we should take\[\pm \sqrt{x ^{2}}?\]

OpenStudy (anonymous):

If you don't, then you only get the positive... \[x^2 = 4 \\ \sqrt{x^2} = \sqrt{4} \\ x = 2\] But x can also be -2, so just only '2' is an incomplete solution

OpenStudy (anonymous):

here after\[\sqrt{x ^{2}}=\sqrt{4}\] \[\left| x \right|=2\] \[x=\pm 2\]

OpenStudy (anonymous):

Ohh, you solved the absolute equation. Okay